OpenStudy (anonymous):

A convex polygon has 10 sides. What is the sum of the measure of the interior angles?

OpenStudy (anonymous):

first find the exterior angle where total exterior angle of a polygon = 360 so then one exterior angle = 360/n = 360/10= 36 then a straight line add up to 180= so interor angle = 180 - 36 =144